OpenStudy (diogop110):

int sen(2y)-sen(y) dy, y=0..1. Please, I'm stuck.

OpenStudy (diogop110):

That's all I could do. I don't see the problem.

OpenStudy (xapproachesinfinity):

what's sen?

OpenStudy (xapproachesinfinity):

you mean sine function here?

OpenStudy (diogop110):

sin, sorry. in portuguese is "sen"

OpenStudy (xapproachesinfinity):

ah okay

OpenStudy (xapproachesinfinity):

why you mentioned y=0...1? Ate any rate do sin(2y)=2sinycosy sin2y-siny=siny(2cosy-1)

OpenStudy (xapproachesinfinity):

then do U-Substitution u=2cosy-1 ===> du=-2sinydy so int (2cosy-1)sinydy=-1/2int (udu)=-1/2 [1/2u^2]+C =-1/4(2cosy-1)^2+C you extend it further

OpenStudy (diogop110):

Thank you @xapproachesinfinity, I was wrong because I was using the values in degrees, but the correct way is in rads. It wasn't matching with Wolfram.

OpenStudy (xapproachesinfinity):

why are you using values anyway? you were asked to find the integral correct? so i don't see the reason to use those values unless the question stated to do so